# Circuit Insights: Towards Interpretability Beyond Activations

### v1

@@ C R P F C R P F C R P F C R P F
WeightLens matches or exceeds activation-based methods Across $\sim$250 features per layer, our weight-based method performs on par with or better than activation maximization methods. It achieves higher scores on Clarity and Responsiveness (Table 1), while activation-based methods tend to overgeneralize, leading to lower scores. However, activation maximization attains higher Purity, highlighting that many features might be context-dependent which is not discovered through WeightLens.
Layer-wise trends in token-based feature interpretability Figure 3 shows that token-based interpretability varies strongly with depth.
Early layers exhibit clear token-level structure, making them well-suited for weight-based analysis, with many features activating reliably on specific tokens. In Gemma-2-2B, however, layers 0 and 7 perform poorly in terms of descriptions quality (Table 1), reflecting their high activation count ($\ell_{0}=76,70$)444huggingface.co/google/gemma-scope-2b-pt-transcoders, which also explains the weak activation-based baselines.
Number of token-based features, as well as the quality of their descriptions, drops sharply in the middle layers of Llama and Gemma, as presented on Figure 3, consistent with prior interpretability analysis (Choi et al., 2024), but not in GPT-2, similartly to the results observed by Bills et al. (2023). A likely explanation is the use of RoPE in Llama and Gemma, which introduces additional non-linearities. Within Gemma-2-2B, layer 12 is the least interpretable based on weights: despite high sparsity ($\ell_{0}=6$), it contains few validated token-based features. Majority of the features in this layer are extremely context-dependent and encode very specific patterns.
Higher layers partially recover in terms of presence of token-based features, though still below early-layer levels. For instance, Gemma-2-2B layer 21 ($\ell_{0}=13$) shows strong interpretability, with token-based features often acting as key–value pairs that map input tokens to predictable collocations (e.g., “apologize for”, “will be”).
Figure 4: Comparison of Neuronpedia feature descriptions for token-based features (i.e., those for which descriptions were successfully generated via WeightLens) and context-based features, whose activations are strongly influenced by context.
Faithfulness is consistently low across layers and methods, likely due to the transcoder architecture: unlike SAEs, which decompose the full residual stream, transcoders write into it like MLPs, so modifying a single feature rarely produces large effects because similar concepts are distributed across layers and features. It may therefore be reasonable to modify the faithfulness metric to evaluate interventions on entire circuits rather than individual features.
Most interpretable features are token-based Only a subset of features receive validated input-invariant descriptions: 32.7% for Gemma, 58.8% for GPT-2, and 25.4% for Llama (Figure 3). However, when the weight-based descriptions fail, activation-based ones also perform poorly (see Figure 4). This is especially visible on layer 21, which overall demonstrated the best results in terms of interpretability.
@@
LLM postprocessing is not essential Although an LLM can produce more general results and filter out noise in the output logits, its utilization is not mandatory in this analysis, since the results are comparable, as seen in Table 1. This is a promising step to reducing reliance on the explainer models.
5 Circuit-Based Interpretability Results
@@
  • CircuitLens-Input (Circuit-based descriptions): descriptions derived from the activation patterns of the feature, obtained via attribution to attention heads;
  • CircuitLens-Full (Circuit-based full descriptions): descriptions based on activation patterns obtained through attributions to attention heads, augmented with tokens, which generation was influenced by the feature.
  • WeightLens (WL) + CircuitLens-Full (Circuit-based full descriptions integrated with WeigthLens results): circuit-based descriptions are enriched with weight-based tokens obtained via WeightLens, which are incorporated when merging cluster-level descriptions into a full feature description.
For generating circuit-based descriptions, we use sampling, as described in the Subsection 3.2, on a relatively small dataset of 24M tokens (see Appendix B). In addition, to eliminate the factor of the dataset influence, we compare the circuit-based analysis performed on the same data, as used in MaxAct*, i.e. on a large dataset with sampling from the top, as presented in Appendix B. These results will be marked by (top), i.e. CircuitLens-Input (top).
As baselines, we consider Neuronpedia and MaxAct* feature desciprions. We do not analyze output-based patterns separately, as they are more computationally expensive to obtain and often provide little additional information without the activation patterns that originally triggered the feature.
Activations alone are not sufficient Figure 5 shows three circuit-based clusters of activating input samples for feature 619 in layer 12 (L12F619). Within each cluster, some commonality in activating concepts exists, but no clear general pattern emerges from activations alone. Isolating the tokens that contribute most through attention heads reveals that each activating entity is either explicitly mentioned or marked as definite or demonstrative references such as “the,” “this,” or “that,” as well as “former” or “latter,” where contributing tokens align semantically.
Extending to the output, feature L21F91 shows that functional roles are not fully captured by input activations: while it activates on tokens like “on” or certain verbs, its main effect is generating output phrases such as “the basis of” or “based on” (Figure 6). These contributions, detectable via attributions from output logits, illustrate how features influence output patterns beyond input activations.
Figure 6: Input patterns for feature L21F91, obtained via attributions to attention heads, correspond to prompts that trigger the feature, and output patterns, obtained via attributions from logits to the feature, are derived from the model’s generated continuations. Tokens isolated through attribution are shown in bold, and activating tokens are highlighted in green.
Downstream effect of a feature Output-based analysis is computationally expensive, as each generated token requires both a forward and backward pass. We generate 15 new tokens per sample to assess how many are needed for reliable results. As expected, early layers (e.g., layers 0 and 7) rarely contribute directly to the output (Figure 7(a)), and when they do, the effect is usually limited to the token immediately following the activating one (Figure 7(b)). Extending the analysis to additional tokens adds only about 1.4% of samples for layer 0, a negligible gain given the computational cost. In deeper layers, most influence remains on the first generated token, though multi-token output patterns also appear, and might be crucial for interpreting a feature’s function, particularly in later layers (e.g., Figure 6).
(a) Smoothed histogram of the fraction of features per layer influencing a given percentage of outputs, with both axes square-root transformed to highlight skewed distributions.
@@
Figure 8: Histogram of number discovered clusters per layer with log scale (y-axis).
Circuit-based polysemanticity Layer 0 has almost no underlying circuit structure, except for its own attention heads, resulting in mainly token-based activations (Figure 3), with few or no clusters beyond outliers. This aligns with its low sparsity, broad topic coverage, and context-independent behavior, yielding on average only 1.05 clusters per feature (see Figure 8). By contrast, Layer 7 exhibits clear polysemanticity, averaging 4.5 clusters per feature and showing the fewest single-cluster cases, indicating widespread circuit formation. Layer 12 presents a mixed picture. While it averages 2.8 clusters per feature, many are either single-cluster or highly clustered. Qualitative inspection suggests that circuit-based clustering can capture both main circuits and sub-circuits, and more careful hyperparameter tuning could yield more generalizable results (Figure 5). It also has the largest share of outlier-only features, reflecting the difficulty of disentangling circuit-driven from semantic-driven activations. Layer 21 is similar (avg. 2.95 clusters/feature) but shows fewer extreme cases. As this layer is closely tied to output generation, we hypothesize that features flexibly participate in multiple circuits to support specific outputs (Figure 6).
Resolving the Clarity Problem As shown by Puri et al. (2025), sparse feature descriptions are often not clear enough, as they fail to specify what precisely triggers a feature’s activation. In Table 2, we compare circuit-based analysis with activation-based baselines, including experiments performed on a smaller dataset sampled from the full distribution, as described in Section 3.2 (CircuitLens-Input and -Full), and on a MaxAct*-like dataset (CircuitLens (top)), which is drawn from a much larger corpus but restricted to top activations (see Appendix B)
Layer 0 Layer 7 Layer 12 Layer 21
C R P F C R P F C R P F C R P F
CircuitLens-Input 0.51 0.20 0.11 0.01 0.39 0.14 0.14 0.03 0.19 0.10 0.10 0.01 0.42 0.59 0.51 0.03
CircuitLens-Full 0.52 0.22 0.11 0.02 0.44 0.14 0.13 0.03 0.24 0.09 0.09 0.01 0.52 0.64 0.54 0.03
CircuitLens-Input (top) 0.66 0.24 0.05 0.02 0.62 0.17 0.08 0.03 0.27 0.09 0.08 0.00 0.54 0.72 0.63 0.03
CircuitLens-Full (top) 0.66 0.24 0.04 0.02 0.61 0.18 0.07 0.04 0.26 0.10 0.07 0.01 0.56 0.73 0.62 0.03
WL + CircuitLens-Full 0.55 0.22 0.10 0.02 0.51 0.14 0.11 0.02 0.25 0.10 0.08 0.01 0.55 0.68 0.56 0.03
Neuronpedia 0.51 0.23 0.09 0.03 0.44 0.17 0.11 0.03 0.15 0.10 0.11 0.00 0.36 0.71 0.64 0.02
MaxAct* 0.50 0.21 0.09 0.04 0.48 0.15 0.09 0.03 0.17 0.10 0.10 0.00 0.39 0.71 0.64 0.02
Table 2: Evaluation of Gemma-2-2b transcoder descriptions (C = Clarity, R = Responsiveness, P = Purity, F = Faithfulness; metrics in range 0–1, higher is better). Best results marked in bold. Methods: Circuit-Based variants. Baselines: Neuronpedia and MaxAct*.
@@ C R P F C R P F C R P F C R P F
Our results show that circuit-based methods still depend on the dataset, with descriptions from the larger dataset achieving the strongest performance across layers, particularly in clarity and responsiveness. However, descriptions derived from the smaller dataset remain competitive and in some cases outperform activation-based baselines, generated on the larger dataset. Combining weight-based and circuit-based analysis further reduces sensitivity to dataset size and distribution, making interpretability more robust. Moreover, analysis of the metrics distributions reveals that circuit-based methods yield far fewer features with extremely low clarity compared to purely activation-based approaches (see Appendix D). Finally, both qualitative and quantitative evidence indicate that sampling from the full distribution, though memory-intensive, provides a more faithful picture of general feature behavior.
@@ Conclusion
In this work, we address a fundamental missing piece in automated interpretability pipelines by developing methods that leverage models’ underlying structural information. We show that raw activations alone often fail to reveal the patterns driving feature activation, while transcoders, well-suited for circuit discovery, enable efficient incorporation of structural information. Our proposed frameworks demonstrate that structural information allow more scalable and robust interpretability. The weight-based analysis offers an efficient alternative for context-independent features – covering up to 58.8% of cases – without requiring large datasets or external LLMs. Circuit-based clustering isolates groups of activating texts into more interpretable clusters, while input- and output-based analysis further clarifies each feature’s functional role. Together, these methods reduce dependence on large datasets, improve robustness, and make automated interpretability more scalable and practical for real-world applications. By bridging activation-based approaches with weight-based analysis and circuit discovery, our work opens new avenues for understanding model behavior at scale.
@@ A.1 Automated Interpretability
Most work in automated interpretability builds on the pipeline of Bills et al. (2023), where a dataset is processed through GPT-2 (Radford et al., 2019) to collect MLP neuron activations. A larger LLM (GPT-4; (OpenAI andAchiam et al., 2024)) then generates descriptions based on top-k activating sequences, and is further used as an “activations simulator” to evaluate these descriptions.
@@ Appendix B Dataset Processing
We used the uncopyrighted version of the Pile dataset Gao et al. (2020), available on Hugging Face (https://huggingface.co/datasets/monology/pile-uncopyrighted) with all copyrighted content removed. This version contains over 345.7 GB of training data from various sources. From this dataset, we extract two datasets of size 6GB (3.6B tokens), and 40MB (24M tokens) for generating MaxAct* descriptions and circuit based descriptions. The extracted portion from the training partition was used to collect the most activated samples based on frequency quantile sampling for the smaller dataset and top percentile sampling for the larger dataset. For evaluations, we utilized the test partition from the same dataset, applying identical preprocessing steps as those used for the training data.

### v2

@@ 4 WeightLens: Evaluations and Findings
Across $\sim$250 features per layer, WeightLens methods perform on par with or better than activation maximization methods. They generally achieve higher scores on Clarity and Responsiveness (Figure 3), while activation-based methods tend to overgeneralize, leading to lower scores. However, activation maximization attains higher Purity, highlighting that many features might be context-dependent, which is not discovered through WeightLens.
@@
Furthermore, these results demonstrate that although an LLM can produce more general results and filter out noise in the output logits, its utilization is not mandatory in this analysis, since the results are comparable.
Figure 5: Description quality gap between token-based and context-dependent features based on Neuronpedia descriptions.
Faithfulness is low across layers and methods, with the rise in the later layers, likely due to the transcoder architecture: in contrast to SAEs, which decompose the full residual stream, transcoders write into it like MLPs. Therefore, steering a single feature rarely produces large effects due to model’s redundancy. Modifying the faithfulness metric to evaluate interventions on groups of features or on entire circuits rather than individual features might bring more trustworthy evaluations.
Layer-wise analysis of the validated features demonstrates, that token-based interpretability varies strongly with depth (see Appendix G for detailed results). Early layers exhibit clear token-level structure, making them well-suited for weight-based analysis, with many features activating reliably on specific tokens. In Gemma-2-2B, however, early layers perform slightly worse in terms of descriptions quality than later ones, as demonstrated in Figure 3, reflecting their high activation count ($\ell_{0}$ of 70–88 for layers 0–7 in contrast to $\ell_{0}$ of 13–41 for layers 18–25 )444huggingface.co/google/gemma-scope-2b-pt-transcoders.
Number of token-based features, as well as the quality of their descriptions, drops in the middle layers of Llama and Gemma, as presented in Figure 4, consistent with prior interpretability analysis (Choi et al., 2024), but not in GPT-2, similarity to the results observed by Bills et al. (2023). A likely explanation is the use of RoPE in Llama and Gemma, which introduces additional non-linearities. Within Gemma-2-2B, layer 12 is the least interpretable based on weights: despite high sparsity ($\ell_{0}=6$), it contains few validated token-based features. Majority of the features in this layer are extremely context-dependent and encode very specific patterns.
Later layers partially recover in terms of presence of token-based features, though still below early-layer levels. For instance, Gemma-2-2B layer 21 ($\ell_{0}=13$) shows strong interpretability, with token-based features often acting as key–value pairs that map input tokens to predictable collocations (e.g., “apologize for”, “will be”).
Additionally, we observe that most interpretable features are token-based. Only a subset of features receive validated input-invariant descriptions: 32.7% for Gemma, 58.8% for GPT-2, and 25.4% for Llama. However, when the weight-based descriptions fail, activation-based ones also perform poorly, as presented in Figure 5.
5 CircuitLens: Evaluations and Findings
@@
  • CircuitLens-Input: descriptions derived from the activation patterns of the feature, obtained via attribution to attention heads;
  • CircuitLens-Full: descriptions based on activation patterns obtained through attributions to attention heads, augmented with tokens, which generation was influenced by the feature.
  • WeightLens (WL) + CircuitLens-Full: circuit-based descriptions are enriched with weight-based tokens obtained via WeightLens, which are incorporated when merging cluster-level descriptions into a full feature description.
Figure 6: Evaluation of Gemma-2-2B transcoder descriptions (C = Clarity, R = Responsiveness, P = Purity, F = Faithfulness). Methods: CircuitLens variants. Baselines: Neuronpedia and MaxAct*.
For generating circuit-based descriptions, we use sampling, as described in the Subsection 3.3, on a relatively small dataset of 24M tokens (see Appendix B). In addition, to eliminate the factor of the dataset influence, we compare the circuit-based analysis performed on the same data, as used in MaxAct*, i.e., on a large dataset (2.3B tokens) with sampling from the top, as presented in Appendix B. These results are marked by (top), e.g., CircuitLens-Input (top). For comparison, we use the same MaxAct* and Neuronpedia baselines as in Section 4. Computational requirements for CircuitLens are demonstrated in Appendix C.
Evaluation results, presented in Figure 6, show that circuit-based methods still depend on the dataset, with descriptions from the larger dataset demonstrating the strongest performance across layers, particularly on input-centric metrics. However, descriptions derived from the smaller dataset remain competitive and in some cases outperform activation-based baselines, generated on the larger dataset. Combining weight-based and circuit-based analysis further reduces sensitivity to dataset size and distribution, making interpretability more robust.
(a) Smoothed histogram of the fraction of features per layer influencing a given percentage of outputs, with both axes square-root transformed.
@@
Sparse feature descriptions often have low Clarity, as demonstrated by Puri et al. (2025): they fail to specify what precisely triggers a feature’s activation. Analysis of the metrics distributions reveals that circuit-based methods yield far fewer features with extremely low clarity compared to purely activation-based approaches (see Appendix G). Additionally, both qualitative and quantitative evidence indicate that sampling from the full distribution, though memory-intensive, provides a more faithful picture of general feature behavior.
Figure 9: Input- and output-based patterns for feature L21F91 obtained via CircuitLens. Tokens isolated through attribution are shown in bold, and activating tokens are highlighted in green.
We demonstrate in Figure 8 that circuit-based clustering and masking of the input allows uncovering patterns that are not visible based on feature’s activation, as illustrated by the feature 619 in layer 12 (L12F619). Within each cluster, some commonality in activating concepts exists, but no clear general pattern emerges from activations alone. Isolating the tokens that contribute most through attention heads reveals that each activating entity is either explicitly mentioned or marked as definite or demonstrative references such as “the” or “this,” as well as “former” or “latter,” where contributing tokens align semantically.
Extending to the output, the analysis of feature L21F91 demonstrates that functional roles are not fully captured by input activations: while it activates on tokens like “on” or certain verbs, its main effect is generating output phrases such as “the basis of” or “based on” (Figure 9). These contributions, detectable via logit attribution, illustrate how features influence the output, generated by the model.
Despite its importance, output-based analysis is computationally expensive, as each generated token requires both a forward and backward pass. We generate 15 new tokens per sample to assess how many are needed for reliable results. As expected, early layers rarely contribute directly to the output, as shown in Figure 7(a), and when they do, the effect is usually limited to the token immediately following the activating one; see Figure 7(b). In deeper layers, most influence remains on the first generated token, though multi-token output patterns also appear, and might be crucial for interpreting a feature’s function, particularly in later layers (e.g., feature L21F91 presented in Figure 9).
Figure 10: Histogram of number discovered clusters per layer with log scale.
Circuit-based clustering aids in interpreting polysemantic features (Appendix E.4), but its effectiveness varies by layer. Layer 0 shows almost no underlying circuit structure beyond its own attention heads, producing mainly token-based activations, according to Figure 4, and averaging only 1.05 clusters per feature, as demonstrated in Figure 10. Layer 7 exhibits clear polysemanticity, with an average of 4.5 clusters per feature and few single-cluster cases, reflecting extensive circuit formation. Layer 12 is mixed, averaging 2.8 clusters per feature, with many single-cluster or highly clustered features. Qualitative inspection suggests circuit-based clustering captures both main circuits and sub-circuits, and hyperparameter tuning could improve generalization, as can be observed in Figure 8. It also has the largest share of outlier-only features, highlighting challenges in disentangling circuit- from semantic-driven activations. Layer 21 is similar (2.95 clusters/feature) but with fewer extreme cases, consistent with its role in output generation and flexible participation of features in multiple circuits, as visible in Figure 9.
Our results demonstrate that both weight-based and circuit-based interpretability provide meaningful insights, improving precision and scalability in automated feature analysis. For future work, CircuitLens could be extended to architectures beyond transcoders, such as SAEs or crosscoders, by leveraging gradient-based attribution methods. Automated clustering of features could be explored further to interpret larger circuits and efficiently identify and steer functional subcomponents. Additionally, more detailed analysis of clustering hyperparameters would help optimize the balance between capturing polysemantic patterns and avoiding overfragmentation, further enhancing the robustness and applicability of automated interpretability across diverse model families.
@@ Conclusion
In this work, we address a fundamental missing piece in automated interpretability pipelines by developing methods that leverage models’ underlying structural information. We show that raw activations alone often fail to reveal the patterns driving feature activation, while transcoders, well-suited for circuit discovery, enable efficient incorporation of structural information. Our proposed frameworks demonstrate that structural information allows more scalable and robust interpretability. The weight-based analysis offers an efficient alternative for context-independent features – covering up to 58.8% of cases – without requiring large datasets or external LLMs. Circuit-based clustering isolates groups of activating texts into more interpretable clusters, while input- and output-based analysis further clarifies each feature’s functional role. Together, these methods reduce dependence on large datasets, improve robustness, and make automated interpretability more scalable and practical for real-world applications. By bridging activation-based approaches with weight-based analysis and circuit discovery, our work opens new avenues for understanding model behavior at scale.
Limitations
WeightLens enables feature interpretation without relying on a dataset or an explainer LLM, but is specific to the transcoder architecture and cannot be directly applied to SAEs. In addition, context-dependent features are not well captured by this approach. CircuitLens improves robustness to dataset size and distribution, and reduces the complexity of the LLM’s task, yet fully eliminating the need for a dataset or an explainer LLM remains challenging. Measuring the downstream effects of features in transcoders is also limited due to high feature redundancy and their architectural differences from residual SAEs that demonstrate good results in steering. As a result, faithfulness scores remain low across all experiments, including the baselines.
Acknowledgments
We thank Patrick Kahardipraja for valuable discussions and support. This work was in part supported by the European Union’s Horizon Europe research and innovation programme (EU Horizon Europe) under grant TEMA (101093003), and by the German Research Foundation (DFG) as research unit DeSBi [KI-FOR 5363] (459422098).
@@ A.1 Automated Interpretability
Most work in automated interpretability builds on the pipeline of Bills et al. (2023), where a dataset is passed through GPT-2 (Radford et al., 2019) to collect MLP neuron activations. A larger LLM (GPT-4; (OpenAI andAchiam et al., 2024)) then generates descriptions based on top-k activating sequences, and is further used as an “activations simulator” to evaluate these descriptions.
@@ A.2 Evaluation metrics
In this work, we use the FADE framework to evaluate the obtained descriptions (Puri et al., 2025). It implements the following metrics:
  • Clarity This metric uses synthetic data generated by an explainer LLM based on the proposed concept description. Feature activations on the synthetic data are compared to a reference distribution of activations on a natural dataset. The Gini coefficient quantifies how well these distributions separate, yielding a score of 1 when the synthetic-data activations are substantially stronger, and 0 when they are indistinguishable. This metric evaluates how clearly the description captures the underlying concept, which is crucial for narrow, fine-grained sparse features.
  • Responsiveness An LLM-as-a-judge rates how strongly samples from the natural dataset relate to the described concept (on a 0–2 scale). The Gini coefficient then measures whether the feature activates more strongly on concept-related samples than on uniformly drawn natural data, producing a score between 0 and 1. This reflects how responsive the feature is to the described concept.
  • Purity Using the same data as Responsiveness, this metric replaces the Gini coefficient with Average Precision to quantify whether the feature activates exclusively on the described concept or also on unrelated phenomena. High purity indicates that the description captures all major activation patterns of the feature.
  • Faithfulness This metric evaluates the downstream influence of the feature. The feature’s activation is perturbed, positively scaled, negatively scaled, or fully ablated, and an LLM judges whether the frequency of the described concept in the model’s output changes accordingly.
@@ Appendix B Dataset Processing
We used the uncopyrighted version of the Pile dataset555huggingface.co/datasets/monology/pile-uncopyrighted (Gao et al., 2020) with all copyrighted content removed. This version contains over 345.7 GB of training data from various sources. From this dataset, we extract two datasets of size 6GB (2.3B tokens), and 40MB (24M tokens) for generating MaxAct* descriptions and circuit based descriptions. The extracted portion from the training partition was used to collect the most activated samples based on frequency quantile sampling for the smaller dataset and top percentile sampling for the larger dataset. For evaluations, we utilized the test partition from the same dataset, applying identical preprocessing steps as those used for the training data.